60 percent of women and 45 percent of men who took a class to quit smoking remained nonsmokers
for at least 1 year after completing the class. Those nonsmokers then attended a success party at the
end of the year. If 60% of the original class was male,
a. What percentage of the people attending the party were women?
b. What percentage of the original class attended the party?
Answer #1

60% of the original class was male, and hence 40% are female.

Let us assume there were 100 people in the class. So, there would be 60 males and 40 females in the class

a. Now, out of the 40 females, 60% who quit smoking remained non smokers. So, (0.6*40)=24 women attended the party.

Out of the 60 males, 45% who quit smoking remained non smokers. So, (0.45*60)=27 men attended the party.

Total number of people who attended the party = 24+27=51

Percentage of the people attending the party were women = \frac{24}{51}*100=47.06

ANSWER: 47.06%

b. We have assumed the total class size to be 100. Out of the 100 people, 51 attended the party.

Percentage of the original class who attended the party = \frac{51}{100}*100=51

ANSWER: 51%